OpenStudy (anonymous):

Can someone help show me how they solve for 2(x – 4) – 12 = 2 – 3(x + 4) ?

OpenStudy (anonymous):

you have to use pemdas (2x-8) - 12 = -1x-4

OpenStudy (anonymous):

@CynthiaJune94

OpenStudy (anonymous):

1. expand the paranthesis 2. rearrange to get all "x" terms on the left and all other terms to the right side of the "=" sign 3. simplify

OpenStudy (anonymous):

2(x-4) -12 = 2 -3(x+4) 2x-8-12 = 2-3x-12 2x-20 = -3x-10 +20 +20 2x = -3x+10 +3x +3x 5x = 10 divide both sides by 5 x=2